# Scaling Behavior of Discrete Diffusion Language Models

### v1

@@ 1 Introduction
Diffusion language models (DLMs) have recently emerged as an alternative to autoregressive language models (ALMs), promising to address some fundamental limitations plaguing ALMs such as the inability to generate multiple tokens in parallel as well as the inability to revise previously generated tokens (li2025survey). While DLMs’ performance at small scales lags behind autoregressive models, they have the potential to solve both of these limitations by decomposing the generative process into a sequence of denoising steps where the entire generated sequence of $N$ tokens is gradually refined, starting at pure noise and transforming it to pure signal over the course of $T$ denoising steps. The freedom to choose $T$ independently of $N$ enables the generation of multiple tokens in each step, while also retaining the ability to update every token at every step.
@@
The likelihood gap between autoregressive, masked diffusion, and uniform diffusion models can be explained, at least in part, through the lens of task difficulty: MDMs are trained to generate the data in any random order, which includes, but is not limited to, generating the data in its natural, autoregressive order and is therefore a strictly more difficult problem (kim2025train). Similarly, uniform diffusion can be understood as a strictly more difficult version of masked diffusion where the model has to predict which tokens are noisy and which are noise-free in addition to subsequently imputing the noisy tokens [^3] To see this, consider a hypothetical uniform diffusion scenario where we are given additional information about which tokens are noisy and which are noise free. In this case, the denoising problem becomes equivalent to masked diffusion as it suffices to fill in the noisy (missing) tokens. (amin2025masking). Put differently, going from autoregression to masking to uniform diffusion imposes progressively less structure on the generative process and therefore provides less inductive bias, suggesting that a more expressive model is required to learn the task effectively. Crucially, the scaling behavior of uniform and hybrid-noise DLMs remains an open question, with existing work being limited to small-scale ablations. Furthermore, prior work on scaling MDMs (nie2025scaling) makes some potentially undesirable design choices, such as assuming that the training loss can approach zero given infinite compute as well as fixing the learning rate and batch size to a constant value.
@@
In this work, we refine the strategy from nie2025scaling by putting additional care on tuning crucial hyperparameters and comparing both compute- and token-bound scaling laws for different noise types, including masked, uniform and hybrid-noise diffusion. Our contributions are three-fold:
@@
Figure 2: Compute-optimal token-to-parameter ratios as a function of model size can vary significantly for different training objectives: While ALMs generally call for more training tokens relative to parameters, as shown by Chinchilla (hoffmann2022training), Llama 3 (grattafiori2024llama), and DeepSeek (bi2024deepseek), discrete diffusion models require comparatively more parameters. For masked diffusion, there is a noticeable disagreement between our results and the literature, with nie2025scaling predicting more parameter-heavy scaling and ni2025training predicting more token-heavy scaling.
@@
(1) Diffusion process. To aid with scaling across different noise types, we propose a new family of hybrid diffusion that allows us to easily and smoothly interpolate between masked and uniform diffusion by defining a transition point from masking to uniform diffusion depending on the signal-to-noise-ratio (SNR). We argue that defining the diffusion process through SNR rather than time is more natural and more principled, having become the standard for continuous-state diffusion (kingma2021variational; kingma2023understanding; karras2024analyzing). To derive the ELBO of the proposed diffusion process, we frame it as an instance of generalized interpolating discrete diffusion (GIDD; von2025generalized) and reparameterize the GIDD ELBO in terms of SNR. This reparameterization simplifies both theory and implementation, while also closing the gap to continuous-state diffusion theory and showing that interpolating discrete diffusion, like continuous diffusion, is invariant to the noise schedule (kingma2021variational).
@@
(2) Methodology. We then systematically analyze the scaling behavior across all noise types (masking, uniform, and hybrid), model sizes, training durations, and batch sizes. To aid with scaling, we utilize CompleteP (dey2025completeP) for stable learning rate transfer across model width and depth. Instead of fixing the batch size to a constant value, as is often done in prior work on scaling laws, we find it to be a crucial hyperparameter with an optimal value depending on the training token budget. Thus, it requires careful tuning at each scale, leading us to estimate the scaling laws without learning rate annealing in order to cope with this additional scaling dimension. This is motivated by the recent trend of treating pre-training and annealing as two distinct training stages conducted on potentially different datasets (swissai2025apertus; allal2025smollm2), as well as our own ablations showing that training with and without annealing yields similar optima and a similar loss (up to some constant factor).
@@ 2.1 Discrete Diffusion Models
Discrete diffusion models are a special application of diffusion models (sohl2015deep) to discrete data. Conceptually, diffusion models are trained to reverse a corruption process that gradually transforms clean data into pure noise, also referred to as the prior distribution. This allows them to generate novel data by gradually removing noise from some noisy sample, starting at pure noise and gradually moving towards cleaner and cleaner version. Diffusion models define a forward noising process $q_{t|s}(z_{t}|z_{s})$ with $0\leq s<t\leq t$, which is a Markov chain that gradually adds noise to the latent variable $z_{s}$, starting at $z_{0}:=x$ and gradually increasing the noise level until eventually, at $t=1$, the prior distribution $p_{1}(z_{1})$ consisting of pure noise is reached. The denoising model $p_{\theta}(z_{s}|z_{t})$ is then trained to remove noise by matching the (conditional) backward denoising process $q_{s|t}(z_{s}|z_{t},x)$. Discrete diffusion models (austin2021d3pm) are a special case of this paradigm where the Markov chain operates on a discrete state spaces. In this case, the transitions and marginals of the Markov chain are categorical distribution. For conciseness, we use a vector-based notation for these categorical distributions, writing ${\bm{q}}_{t}(x)$ and ${\bm{q}}_{t|s}(z_{s})$ to denote the marginals $q_{t}(\cdot|x)$ and the Markov transitions $q_{t|s}(\cdot|z_{s})$ respectively.
@@ 2.2 Generalized Interpolating Discrete Diffusion
We adopt generalized interpolating discrete diffusion (GIDD; von2025generalized), a class of discrete diffusion models (austin2021d3pm) that provides a unified perspective of many existing approaches such as masked diffusion (ou2025your; sahoo2024simple; shi2024simplified) or uniform diffusion (schiff2024simple; sahoo2025diffusion). GIDD defines the noising process to be an interpolated categorical distribution between some initial state $z_{0}:=x\in\mathcal{X}$ (the data) and some arbitrary (smoothly time-varying) mixing distribution $\bm{\pi}_{t}$ over latent (noisy) variables $z_{t\in(0,1]}\in\mathcal{Z}\supseteq\mathcal{X}$. Specifically, for some mixing rate $\alpha_{t}$ which determines the signal strength over time, the Markov chain transitions and marginal distributions are given by
@@
We adopt the framework of von2025generalized as it allows us to train discrete diffusion models with different noising properties within a shared framework, reducing precisely to specialized variants in the literature under an appropriate mixing schedule. However, we improve this framework by showing how it can be reformulated in terms of signal-to-noise ratio (SNR), obtaining a simpler, more flexible likelihood bound and closing the gap to continuous-state diffusion theory.
@@ 2.3 Reframing GIDD in Terms of SNR
It is well-known that continuous-state diffusion models are invariant to the noise schedule (kingma2021variational), with many approaches relying on this fact to accelerate training via adaptive noise schedules (kingma2023understanding; karras2024analyzing; dieleman2024schedules). This stems from the insight that the notion of time in diffusion models is spurious and serves only as a proxy for the signal-to-noise ratio (SNR), and that SNR is sufficient and, arguably, a more natural way to describe the forward and backward diffusion process. Similar results have been shown for the special case of masked diffusion (shi2024simplified; sahoo2024simple). In this section, we show that this invariance continues to hold for general interpolating discrete diffusion models following the proof technique by kingma2021variational.
@@ 2.4 A Universal Hybrid Mixing Distribution
Hybrid-noise diffusion (gu2022vector; von2025generalized; haxholli2025efficient) has been proposed as a way to equip masked diffusion models with the ability to revise tokens throughout the denoising process while having a smaller likelihood gap compared to fully uniform diffusion. For our scaling experiments, we consider a mixing distribution $\bm{\pi}_{\lambda}$ that smoothly transitions from masked to uniform diffusion, covering a range of hybrid mixtures in between. Our idea is to interpolate between the pure masking and pure uniform noise based on the log-SNR $\lambda$, thereby controlling how much masking and how much random perturbation happens proportionally at any point of the noising process. We define
@@ 2.5 Anisotropic Noise and Diffusion Forcing
While sequence diffusion models typically use a global isotropic noise level that is shared by all tokens in the sequence, Diffusion Forcing (chen2024diffusion) proposes to sample noise levels independent for each tokens, resulting in anisotropic noise. This is done to stabilize autoregressive rollouts and generally provides enhanced flexibility at inference time as it effectively serves as an augmentation over noise levels. This idea has since been extended to DLMs (wang2025diffusion) to speed up inference.
@@ 3 Estimating scaling laws
Scaling laws have become an important ingredient of large-scale neural network training, particularly in the context of training LLMs. Due to the vast costs associated with large-scale training runs, key decisions are based on forecasts obtained through extrapolating the performance of smaller runs to the desired, bigger scale. Prior work on the scaling of MDMs (nie2025scaling) has made some assumptions that we would like to revisit. For example, the learning rate and batch size are fixed to constant values across all experiments, but this may not be optimal for different model sizes and token budgets (bergsma2025power). Additionally, the reported scaling law is the result of a power law fit without constant offset, thereby implicitly assuming that the ideal training loss is zero and can be reached given infinite compute, which is known not to be the case for ALMs. These limitations prompt us to rederive the scaling laws from scratch, dropping any assumptions on the optimal batch size, learning rate, and irreducible loss. Our scaling laws are estimated directly on the negative ELBO, [^5] In a slight abuse of terminology, we sometimes refer to the negative ELBO simply as the “ELBO”. which constitutes an upper bound on the negative log-likelihood (NLL). While our recipe largely follows the methodology by (hoffmann2022training), which is well-established and has been widely adopted for estimating scaling laws (touvron2023llama; bi2024deepseek; shuai2024scaling), there are some key differences.
@@
While adopting CompleteP enables learning rate transfer across model scales, the same learning rate is not optimal for different batch sizes and training horizons. It is therefore still necessary to sweep the learning rate for each model and batch size in order to find the compute-optimal Pareto frontier. To cope with the computational demands of sweeping the batch size in addition to model and data size, we omit learning rate annealing and analyze the scaling behavior without it. This allows capturing all possible training horizons in a single run per model, batch size, and learning rate. This decision is justified twofold: First, modern large-scale training often treats the annealing phase as distinct from pre-training where the data mixture is often adapted to more closely resemble the test distribution by injecting more high-quality data geared towards the desired downstream tasks (swissai2025apertus; allal2025smollm2). Second, we conduct small-scale ablations to study the effect of omitting annealing, finding that optimal hyperparameters are preserved and that the performance difference is a constant factor (see Sec. 4.4).
@@
Sweeping the learning rate across batch and model sizes reveals the clear existence of a compute- and token-optimal batch size that scales almost linearly in the number of training tokens (Fig. 3). Similar findings have been reported for ALMs when training below the critical batch size (hu2024minicpm; shuai2024scaling; bergsma2025power). The critical batch size refers to the phenomenon where scaling the batch size past a certain critical point yields diminishing returns and becomes compute-inefficient. It is worth noting that we find no dependence of the optimal batch size on the target loss, a claim that has been raised for the critical batch size of ALMs (zhang2024does) and also more generally (mccandlish2018empirical). As our experiments show no signs of saturation even at batch sizes at $10^{6}$ tokens, this suggests that the critical batch size of DLMs lies well above that of ALMs, which has been reported to saturate around $10^{6}$ tokens (shuai2024scaling; zhang2024does). Another hyperparameter that is known to have optimal values depending on the batch size is Adam’s $\beta_{2}$ parameter. However, also for this parameter we find little benefit in deviating from our default value of $0.99$: Neither do we observe benefits from using larger values for small batch sizes [^6] This finding is particular to half-precision training in bfloat16, as we did observe slight benefits from increasing $\beta_{2}$ for full-precision training at low batch sizes. nor from using smaller values at larger batch sizes. [^7] This improved consistency may be a result of using the LaProp (ziyin2020laprop) variant of Adam, or due to the CompleteP (dey2025completeP) parameterization, although we do not investigate this further. We do decrease $\beta_{2}$ to $0.98$ starting at batch sizes of $256$ as we found this to slightly improve stability without any noticeable performance degradation.
@@ 4.1 Model Architecture and Training
Architecture. Our model architecture follows a standard Transformer (vaswani2017attention) with some key modifications. As described in Section 3, we implement CompleteP (dey2025completeP) for optimal learning rate transfer across width and depth. We use Squared ReLU for MLP activations, as recommended by so2021searching. To ensure stable training, we add RMSNorm (zhang2019root) layers without bias before each attention and MLP block following LLaMA (touvron2023llama) as well as to both keys and queries, following QK-norm (naseer2021intriguing; dehghani2023scaling). In the same spirit, we also employ attention logit soft-capping (gemma2024gemma2). Finally, we add attention sinks in the form of attention biases (sun2024massive) to further stabilize training and prevent outlier features (sun2024massive; he2024understanding).
Data. We use Nemotron-CC (su2024nemotron) without quality filtering as a representative dataset of internet-scale pre-training. Since it is known that a larger vocabulary facilitates better scaling (takase2024large; huang2025over) and to ensure efficient tokenization, we train a BPE tokenizer (gage1994bpe; sennrich2015neural) with a vocabulary size of $2^{17}$ (131,072) tokens on a 256 GB subset of the data. The trained tokenizer is released with the model.
@@
We design the diffusion process by aiming to maximize flexibility of the resulting model at inference time, supporting conditional prompt completion, advanced sampling algorithms, as well as flexible length generation. For conditional prompt completion, we select 20% of samples and leave the first $N\cdot\arccos(r)]$, $r\sim\mathcal{U}(0,1)$ tokens noise-free. Attention from prompt queries to completion keys is masked in order to enable KV-caching of the prompt during inference. To support both isotropic and anisotropic denoising, we implement diffusion forcing (chen2024diffusion) by sampling independent per-token noise levels for 50% of samples. Finally, we augment the context with a random fraction $f\sim\mathcal{U}(0,0.2)$ of empty tokens following Dreamon2025 to add some flexibility to the length of generated samples.
Optimization. Instead of directly minimizing the ELBO, we use the unweighted ELBO (Eq. 5 with $p(\lambda):=1$) as a surrogate loss, as this has been found to give better convergence for both hybrid and masked diffusion models (von2025generalized; sahoo2025esoteric). Following hafner2023mastering, we use LaProp (ziyin2020laprop) over Adam for its improved stability on a wider range of $\beta_{2}$ and $\epsilon$ values. The learning rate is warmed up over the first 2000 steps of training and held constant, with most experiments not including a cooldown phase. For the experiments that do have a cooldown phase, we anneal the learning rate to 0 over the last 20% of training following the WSD schedule (hu2024minicpm; hagele2024scaling).

### v2

@@ 1 Introduction
Diffusion language models (DLMs) have recently emerged as an alternative to autoregressive language models (ALMs), promising to address some fundamental limitations plaguing ALMs such as the inability to generate multiple tokens in parallel as well as the inability to revise previously generated tokens (Li et al., 2025). While DLMs’ performance at small scales lags behind autoregressive models, they have the potential to solve both of these limitations by decomposing the generative process into a sequence of denoising steps where the entire generated sequence of $N$ tokens is gradually refined, starting at pure noise and transforming it to pure signal over the course of $T$ denoising steps. The freedom to choose $T$ independently of $N$ enables the generation of multiple tokens in each step, while also retaining the ability to update every token at every step.
@@
The likelihood gap between autoregressive, masked diffusion, and uniform diffusion models can be explained, at least in part, through the lens of task difficulty: MDMs are trained to generate the data in any random order, which includes, but is not limited to, generating the data in its natural, autoregressive order and is therefore a strictly more difficult problem (Kim et al., 2025). Similarly, uniform diffusion can be understood as a strictly more difficult version of masked diffusion where the model has to predict which tokens are noisy and which are noise-free in addition to subsequently imputing the noisy tokens [^3] To see this, consider a hypothetical uniform diffusion scenario where we are given additional information about which tokens are noisy and which are noise free. In this case, the denoising problem becomes equivalent to masked diffusion as it suffices to fill in the noisy (missing) tokens. (Amin et al., 2025). Put differently, going from autoregression to masking to uniform diffusion imposes progressively less structure on the generative process and therefore provides less inductive bias, suggesting that a more expressive model is required to learn the task effectively. Crucially, the scaling behavior of uniform and hybrid-noise DLMs remains an open question, with existing work being limited to small-scale ablations. Furthermore, prior work on scaling MDMs (Nie et al., 2025a) makes some potentially undesirable design choices, such as assuming that the training loss can approach zero given infinite compute as well as fixing the learning rate and batch size to a constant value.
@@
In this work, we refine the strategy from Nie et al. (2025a) by putting additional care on tuning crucial hyperparameters and comparing both compute- and token-bound scaling laws for different noise types, including masked, uniform and hybrid-noise diffusion. Our contributions are three-fold:
@@
Figure 2: Compute-optimal token-to-parameter ratios as a function of model size can vary significantly for different training objectives: While ALMs generally call for more training tokens relative to parameters, as shown by Chinchilla (Hoffmann et al., 2022), Llama 3 (Grattafiori et al., 2024), and DeepSeek (Bi et al., 2024), discrete diffusion models require comparatively more parameters. For masked diffusion, there is a noticeable disagreement between our results and the literature, with Nie et al. (2025a) predicting more parameter-heavy scaling and Ni et al. (2025) predicting more token-heavy scaling.
@@
(1) Diffusion process. To aid with scaling across different noise types, we propose a new family of hybrid diffusion that allows us to easily and smoothly interpolate between masked and uniform diffusion by defining a transition point from masking to uniform diffusion depending on the signal-to-noise-ratio (SNR). We argue that defining the diffusion process through SNR rather than time is more natural and more principled, having become the standard for continuous-state diffusion (Kingma et al., 2021; Kingma & Gao, 2023; Karras et al., 2024). To derive the ELBO of the proposed diffusion process, we frame it as an instance of generalized interpolating discrete diffusion (GIDD; von Rütte et al., 2025) and reparameterize the GIDD ELBO in terms of SNR. This reparameterization simplifies both theory and implementation, while also closing the gap to continuous-state diffusion theory and showing that interpolating discrete diffusion, like continuous diffusion, is invariant to the noise schedule (Kingma et al., 2021).
@@
(2) Methodology. We then systematically analyze the scaling behavior across all noise types (masking, uniform, and hybrid), model sizes, training durations, and batch sizes. To aid with scaling, we utilize CompleteP (Dey et al., 2025) for stable learning rate transfer across model width and depth. Instead of fixing the batch size to a constant value, as is often done in prior work on scaling laws, we find it to be a crucial hyperparameter with an optimal value depending on the training token budget. Thus, it requires careful tuning at each scale, leading us to estimate the scaling laws without learning rate annealing in order to cope with this additional scaling dimension. This is motivated by the recent trend of treating pre-training and annealing as two distinct training stages conducted on potentially different datasets (Project Apertus, 2025; Allal et al., 2025), as well as our own ablations showing that training with and without annealing yields similar optima and a similar loss (up to some constant factor).
@@ 2.1 Discrete Diffusion Models
Discrete diffusion models are a special application of diffusion models (Sohl-Dickstein et al., 2015) to discrete data. Conceptually, diffusion models are trained to reverse a corruption process that gradually transforms clean data into pure noise, also referred to as the prior distribution. This allows them to generate novel data by gradually removing noise from some noisy sample, starting at pure noise and gradually moving towards cleaner and cleaner version. Diffusion models define a forward noising process $q_{t|s}(z_{t}|z_{s})$ with $0\leq s<t\leq t$, which is a Markov chain that gradually adds noise to the latent variable $z_{s}$, starting at $z_{0}:=x$ and gradually increasing the noise level until eventually, at $t=1$, the prior distribution $p_{1}(z_{1})$ consisting of pure noise is reached. The denoising model $p_{\theta}(z_{s}|z_{t})$ is then trained to remove noise by matching the (conditional) backward denoising process $q_{s|t}(z_{s}|z_{t},x)$. Discrete diffusion models (Austin et al., 2021) are a special case of this paradigm where the Markov chain operates on a discrete state spaces. In this case, the transitions and marginals of the Markov chain are categorical distribution. For conciseness, we use a vector-based notation for these categorical distributions, writing ${\bm{q}}_{t}(x)$ and ${\bm{q}}_{t|s}(z_{s})$ to denote the marginals $q_{t}(\cdot|x)$ and the Markov transitions $q_{t|s}(\cdot|z_{s})$ respectively.
@@ 2.2 Generalized Interpolating Discrete Diffusion
We adopt generalized interpolating discrete diffusion (GIDD; von Rütte et al., 2025), a class of discrete diffusion models (Austin et al., 2021) that provides a unified perspective of many existing approaches such as masked diffusion (Ou et al., 2025; Sahoo et al., 2024; Shi et al., 2024) or uniform diffusion (Schiff et al., 2024; Sahoo et al., 2025a). GIDD defines the noising process to be an interpolated categorical distribution between some initial state $z_{0}:=x\in\mathcal{X}$ (the data) and some arbitrary (smoothly time-varying) mixing distribution $\bm{\pi}_{t}$ over latent (noisy) variables $z_{t\in(0,1]}\in\mathcal{Z}\supseteq\mathcal{X}$. Specifically, for some mixing rate $\alpha_{t}$ which determines the signal strength over time, the Markov chain transitions and marginal distributions are given by
@@
We adopt the framework of von Rütte et al. (2025) as it allows us to train discrete diffusion models with different noising properties within a shared framework, reducing precisely to specialized variants in the literature under an appropriate mixing schedule. However, we improve this framework by showing how it can be reformulated in terms of signal-to-noise ratio (SNR), obtaining a simpler, more flexible likelihood bound and closing the gap to continuous-state diffusion theory.
@@ 2.3 Reframing GIDD in Terms of SNR
It is well-known that continuous-state diffusion models are invariant to the noise schedule (Kingma et al., 2021), with many approaches relying on this fact to accelerate training via adaptive noise schedules (Kingma & Gao, 2023; Karras et al., 2024; Dieleman, 2024). This stems from the insight that the notion of time in diffusion models is spurious and serves only as a proxy for the signal-to-noise ratio (SNR), and that SNR is sufficient and, arguably, a more natural way to describe the forward and backward diffusion process. Similar results have been shown for the special case of masked diffusion (Shi et al., 2024; Sahoo et al., 2024). In this section, we show that this invariance continues to hold for general interpolating discrete diffusion models following the proof technique by Kingma et al. (2021).
@@ 2.4 A Universal Hybrid Mixing Distribution
Hybrid-noise diffusion (Gu et al., 2022; von Rütte et al., 2025; Haxholli et al., 2025) has been proposed as a way to equip masked diffusion models with the ability to revise tokens throughout the denoising process while having a smaller likelihood gap compared to fully uniform diffusion. For our scaling experiments, we consider a mixing distribution $\bm{\pi}_{\lambda}$ that smoothly transitions from masked to uniform diffusion, covering a range of hybrid mixtures in between. Our idea is to interpolate between the pure masking and pure uniform noise based on the log-SNR $\lambda$, thereby controlling how much masking and how much random perturbation happens proportionally at any point of the noising process. We define
@@ 2.5 Anisotropic Noise and Diffusion Forcing
While sequence diffusion models typically use a global isotropic noise level that is shared by all tokens in the sequence, Diffusion Forcing (Chen et al., 2024) proposes to sample noise levels independent for each tokens, resulting in anisotropic noise. This is done to stabilize autoregressive rollouts and generally provides enhanced flexibility at inference time as it effectively serves as an augmentation over noise levels. This idea has since been extended to DLMs (Wang et al., 2025) to speed up inference.
@@ 3 Estimating scaling laws
Scaling laws have become an important ingredient of large-scale neural network training, particularly in the context of training LLMs. Due to the vast costs associated with large-scale training runs, key decisions are based on forecasts obtained through extrapolating the performance of smaller runs to the desired, bigger scale. Prior work on the scaling of MDMs (Nie et al., 2025a) has made some assumptions that we would like to revisit. For example, the learning rate and batch size are fixed to constant values across all experiments, but this may not be optimal for different model sizes and token budgets (Bergsma et al., 2025). Additionally, the reported scaling law is the result of a power law fit without constant offset, thereby implicitly assuming that the ideal training loss is zero and can be reached given infinite compute, which is known not to be the case for ALMs. These limitations prompt us to rederive the scaling laws from scratch, dropping any assumptions on the optimal batch size, learning rate, and irreducible loss. Our scaling laws are estimated directly on the negative ELBO, [^5] In a slight abuse of terminology, we sometimes refer to the negative ELBO simply as the “ELBO”. which constitutes an upper bound on the negative log-likelihood (NLL). While our recipe largely follows the methodology by (Hoffmann et al., 2022), which is well-established and has been widely adopted for estimating scaling laws (Touvron et al., 2023; Bi et al., 2024; Shuai et al., 2024), there are some key differences.
@@
While adopting CompleteP enables learning rate transfer across model scales, the same learning rate is not optimal for different batch sizes and training horizons. It is therefore still necessary to sweep the learning rate for each model and batch size in order to find the compute-optimal Pareto frontier. To cope with the computational demands of sweeping the batch size in addition to model and data size, we omit learning rate annealing and analyze the scaling behavior without it. This allows capturing all possible training horizons in a single run per model, batch size, and learning rate. This decision is justified twofold: First, modern large-scale training often treats the annealing phase as distinct from pre-training where the data mixture is often adapted to more closely resemble the test distribution by injecting more high-quality data geared towards the desired downstream tasks (Project Apertus, 2025; Allal et al., 2025). Second, we conduct small-scale ablations to study the effect of omitting annealing, finding that optimal hyperparameters are preserved and that the performance difference is a constant factor (see Sec. 4.4).
@@
Sweeping the learning rate across batch and model sizes reveals the clear existence of a compute- and token-optimal batch size that scales almost linearly in the number of training tokens (Fig. 3). Similar findings have been reported for ALMs when training below the critical batch size (Hu et al., 2024; Shuai et al., 2024; Bergsma et al., 2025). The critical batch size refers to the phenomenon where scaling the batch size past a certain critical point yields diminishing returns and becomes compute-inefficient. It is worth noting that we find no dependence of the optimal batch size on the target loss, a claim that has been raised for the critical batch size of ALMs (Zhang et al., 2024) and also more generally (McCandlish et al., 2018). As our experiments show no signs of saturation even at batch sizes at $10^{6}$ tokens, this suggests that the critical batch size of DLMs lies well above that of ALMs, which has been reported to saturate around $10^{6}$ tokens (Shuai et al., 2024; Zhang et al., 2024). Another hyperparameter that is known to have optimal values depending on the batch size is Adam’s $\beta_{2}$ parameter. However, also for this parameter we find little benefit in deviating from our default value of $0.99$: Neither do we observe benefits from using larger values for small batch sizes [^6] This finding is particular to half-precision training in bfloat16, as we did observe slight benefits from increasing $\beta_{2}$ for full-precision training at low batch sizes. nor from using smaller values at larger batch sizes. [^7] This improved consistency may be a result of using the LaProp (Ziyin et al., 2020) variant of Adam, or due to the CompleteP (Dey et al., 2025) parameterization, although we do not investigate this further. We do decrease $\beta_{2}$ to $0.98$ starting at batch sizes of $256$ as we found this to slightly improve stability without any noticeable performance degradation.
@@ 4.1 Model Architecture and Training
Architecture. Our model architecture follows a standard Transformer (Vaswani et al., 2017) with some key modifications. As described in Section 3, we implement CompleteP (Dey et al., 2025) for optimal learning rate transfer across width and depth. We use Squared ReLU for MLP activations, as recommended by So et al. (2021). To ensure stable training, we add RMSNorm (Zhang & Sennrich, 2019) layers without bias before each attention and MLP block following LLaMA (Touvron et al., 2023) as well as to both keys and queries, following QK-norm (Naseer et al., 2021; Dehghani et al., 2023). In the same spirit, we also employ attention logit soft-capping (Gemma Team, 2024). Finally, we add attention sinks in the form of attention biases (Sun et al., 2024) to further stabilize training and prevent outlier features (Sun et al., 2024; He et al., 2024).
Data. We use Nemotron-CC (Su et al., 2024) without quality filtering as a representative dataset of internet-scale pre-training. Since it is known that a larger vocabulary facilitates better scaling (Takase et al., 2024; Huang et al., 2025) and to ensure efficient tokenization, we train a BPE tokenizer (Gage, 1994; Sennrich et al., 2015) with a vocabulary size of $2^{17}$ (131,072) tokens on a 256 GB subset of the data. The trained tokenizer is released with the model.
@@
We design the diffusion process by aiming to maximize flexibility of the resulting model at inference time, supporting conditional prompt completion, advanced sampling algorithms, as well as flexible length generation. For conditional prompt completion, we select 20% of samples and leave the first $N\cdot\arccos(r)]$, $r\sim\mathcal{U}(0,1)$ tokens noise-free. Attention from prompt queries to completion keys is masked in order to enable KV-caching of the prompt during inference. To support both isotropic and anisotropic denoising, we implement diffusion forcing (Chen et al., 2024) by sampling independent per-token noise levels for 50% of samples. Finally, we augment the context with a random fraction $f\sim\mathcal{U}(0,0.2)$ of empty tokens following Wu et al. (2025) to add some flexibility to the length of generated samples.
Optimization. Instead of directly minimizing the ELBO, we use the unweighted ELBO (Eq. 5 with $p(\lambda):=1$) as a surrogate loss, as this has been found to give better convergence for both hybrid and masked diffusion models (von Rütte et al., 2025; Sahoo et al., 2025b). Following Hafner et al. (2023), we use LaProp (Ziyin et al., 2020) over Adam for its improved stability on a wider range of $\beta_{2}$ and $\epsilon$ values. The learning rate is warmed up over the first 2000 steps of training and held constant, with most experiments not including a cooldown phase. For the experiments that do have a cooldown phase, we anneal the learning rate to 0 over the last 20% of training following the WSD schedule (Hu et al., 2024; Hägele et al., 2024).